Please explain and show work.Find the maximum and minimum values of 1f(xy)= XY. Subect to the constraint x2-y =12 Assume that y<0 for this problemy
Answer #1

* Given function is f(x,y) = xy. subject to the Constant x²y =12 1 y=x2-12 f(x) = x(x?-12) = x-122 f has maximum cos minimumBy second derivative West, f has minimum values minimum value = f(2)= (2²_12(2) = 8-24 = -16. -12 CO f(-2) = 6(-2) = f has m